An electrical load operates at 240 V rms. The load absorbs an average of 5 KW at lagging power factor of 0.8 The reactive power Q is:


 Select one

 a 3.75 KVAR

 b 4.50 KVAR

 c 6.00 KVAR

 d 6.75 KVAR

 e 5.25 KVAR


1
Expert's answer
2021-06-03T16:22:02-0400

Average power, P "= 5KW"

But "P=/S/ cos \theta"

Given power factor "cos \theta=0.8"

"=sin \theta=\sqrt{1-cos^2\theta}"

"=sin \theta=\sqrt{1-0.8^2}=\sqrt{0.36=0.6}"

"P=/S/ cos \theta"

"5KW=\/S\/*0.8"

"/S/=\frac{5000}{0.8}=6250"

Reactive power, Q "=/S/ sin\theta" VAR

"Q=6250(0.6)=3750" VAR

"S=P+jQ"

"S=5000+j3750"

"S=5+j3.75" KVA